# Con-ReCall: Detecting Pre-training Data in LLMs via Contrastive Decoding

### v1

@@ Abstract
The training data in large language models is key to their success, but it also presents privacy and security risks, as it may contain sensitive information. Detecting pre-training data is crucial for mitigating these concerns. Existing methods typically analyze target text in isolation or solely with non-member contexts, overlooking potential insights from simultaneously considering both member and non-member contexts. While previous work suggested that member contexts provide little information due to the minor distributional shift they induce, our analysis reveals that these subtle shifts can be effectively leveraged when contrasted with non-member contexts. In this paper, we propose Con-ReCall, a novel approach that leverages the asymmetric distributional shifts induced by member and non-member contexts through contrastive decoding, amplifying subtle differences to enhance membership inference. Extensive empirical evaluations demonstrate that Con-ReCall achieves state-of-the-art performance on the WikiMIA benchmark and is robust against various text manipulation techniques.
@@
Cheng Wang, Yiwei Wang§, Bryan Hooi, Yujun Cai, Nanyun Peng§, Kai-Wei Chang§ National University of Singapore § University of California, Los Angeles   Nanyang Technological University wcheng@comp.nus.edu.sg https://github.com/WangCheng0116/CON-RECALL
@@ 1 Introduction
We summarize our contributions as follows: 1) We introduce Con-ReCall, a novel contrastive decoding approach that effectively utilizes both member and non-member contexts, significantly enhancing the distinction between member and non-member data in LLMs. 2) Through extensive experiments, we demonstrate that Con-ReCall achieves substantial improvements over existing baselines, highlighting its effectiveness and resilience in detecting pre-training data. 3) We demonstrate that Con-ReCall is robust against text manipulation techniques, including random deletion, synonym substitution, and paraphrasing, maintaining superior performance and resilience to potential evasion strategies.
@@ 2 Related Work
Membership inference attack.
Membership inference attack (MIA) was first proposed by Shokri et al. (2017). MIA has been extensively studied, particularly in classification models within the computer vision domain (Carlini et al., 2023a; Zarifzadeh et al., 2024; Bertran et al., 2023). While there is growing attention to MIA in language models, most work has focused on detecting fine-tuning data (Watson et al., 2022; Mireshghallah et al., 2022; Fu et al., 2024). MIA can serve as a powerful tool for detecting copyrighted materials (Meeus et al., 2023; Duarte et al., 2024; Chang et al., 2023), personally identifiable information (Mozes et al., 2023; Tang et al., 2024) and test-set contamination (Sainz et al., 2023; Oren et al., 2023).
@@
Although detecting pre-training data is an instance of MIA, it faces greater challenges compared to traditional MIA. Classical MIA (Shokri et al., 2017) typically requires training a shadow model using data sampled from the training data distribution. However, for large language models, many developers are reluctant to release the full training data (OpenAI, 2024a; Touvron et al., 2023b), making it impractical to train shadow models. Additionally, due to the sheer volume of training data, LLMs are usually trained for a single epoch, which makes memorization inherently difficult and detection even more challenging (Carlini et al., 2023b; Shi et al., 2024a).
To our knowledge, Shi et al. (2024a) was the first to investigate this problem, contributing a baseline method and the WikiMIA benchmark. Their method, Min-K%, despite its simplicity, serves as a powerful baseline. Zhang et al. (2024) enhanced Min-Kby normalizing token log-probabilities. The ReCall method (Xie et al., 2024) introduces relative conditional log-likelihoods and achieves current state-of-the-art performance.
@@ 5 Analysis
To further evaluate the effectiveness and practicality of Con-ReCall, we conducted additional analyses focusing on its robustness and adaptability in real-world scenarios. These investigations provide deeper insights into the method’s performance under various challenging conditions.
@@ 5.1 Robustness of Con-ReCall
As membership inference attacks gain prominence, it is crucial to evaluate the robustness of these methods against potential evasion techniques. In real-world scenarios, data may not always be presented in its original form due to various factors such as text preprocessing, natural language variations, or intentional obfuscation. Therefore, a robust membership inference method should maintain its effectiveness even when faced with altered versions of the target data.
To assess the robustness of Con-ReCall, we employ three text manipulation techniques. First, we use Random Deletion, where we randomly remove a certain percentage of words from the original text, using deletion rates of 10%, 15%, and 20% in our experiments. Second, we apply Synonym Substitution, replacing a portion of the words in the text with their synonyms. For this technique, we use substitution rates of 10%, 15%, and 20%, utilizing WordNet (Miller, 1994) for synonym selection. Lastly, we leverage the WikiMIA-paraphrased dataset (Zhang et al., 2024), which offers paraphrased versions of the original WikiMIA Shi et al. (2024a) texts. This dataset, created using ChatGPT [^1] OpenAI. https://chat.openai.com/chat to rephrase the original text while preserving its meaning, provides a standardized benchmark for evaluating robustness against paraphrasing.
We evaluate the effectiveness of baselines and Con-ReCall after transforming texts using the above techniques. Our experiments are conducted using Pythia-6.9B (Biderman et al., 2023) and LLaMA-30B (Touvron et al., 2023a) models on the WikiMIA-32 (Shi et al., 2024a) dataset. Table 3 presents the AUC performance for each method under various text manipulation scenarios. The results demonstrate that Con-ReCall consistently outperforms baseline methods across all text manipulation techniques, maintaining its superior performance even when faced with altered versions of the target data. Notably, Con-ReCall shows particular resilience to synonym substitution and paraphrasing, where it experiences minimal performance degradation compared to other methods. This robustness underscores Con-ReCall’s effectiveness in real-world scenarios where data may undergo various transformations.

### v2

@@ Abstract
The training data in large language models is key to their success, but it also presents privacy and security risks, as it may contain sensitive information. Detecting pre-training data is crucial for mitigating these concerns. Existing methods typically analyze target text in isolation or solely with non-member contexts, overlooking potential insights from simultaneously considering both member and non-member contexts. While previous work suggested that member contexts provide little information due to the minor distributional shift they induce, our analysis reveals that these subtle shifts can be effectively leveraged when contrasted with non-member contexts. In this paper, we propose Con-ReCall, a novel approach that leverages the asymmetric distributional shifts induced by member and non-member contexts through contrastive decoding, amplifying subtle differences to enhance membership inference. Extensive empirical evaluations demonstrate that Con-ReCall achieves state-of-the-art performance on the WikiMIA benchmark and is robust against various text manipulation techniques.  [^1] Our code is available at the following repo: https://github.com/WangCheng0116/CON-RECALL
@@
Cheng Wang, Yiwei Wang|| , Bryan Hooi, Yujun Cai, Nanyun Peng§, Kai-Wei Chang§ National University of Singapore || University of California, Merced § University of California, Los Angeles   University of Queensland wcheng@comp.nus.edu.sg
@@ 2 Related Work
While membership inference attacks (MIA) have been extensively studied in various domains (Shokri et al., 2017; Carlini et al., 2023a; Watson et al., 2022), detecting pre-training data in LLMs presents unique challenges. Unlike classical MIA, LLM developers rarely release full training data (OpenAI, 2024a; Touvron et al., 2023b), and single-epoch training on vast datasets makes memorization detection difficult (Carlini et al., 2023b; Shi et al., 2024a). Shi et al. (2024a) pioneered this research with the WikiMIA benchmark and Min-Kbaseline method. Zhang et al. (2024) improved Min-Kthrough token log-probability normalization, while the ReCall method (Xie et al., 2024) currently achieves state-of-the-art performance using relative conditional log-likelihoods. These methods contribute to the broader application of MIA in detecting copyrighted materials, personally identifiable information, and test-set contamination (Meeus et al., 2023; Mozes et al., 2023; Sainz et al., 2023).
@@ 5 Analysis
To further evaluate the effectiveness and practicality of Con-ReCall, we conducted additional analyses focusing on its robustness and adaptability in real-world scenarios.
@@ 5.1 Robustness of Con-ReCall
As membership inference attacks gain prominence, evaluating their robustness against potential evasion techniques becomes crucial. Real-world data may be altered due to preprocessing, language variations, or intentional obfuscation. Therefore, a robust membership inference method should remain effective when faced with modified target data. To assess Con-ReCall’s robustness, we employ three text manipulation techniques. First, Random Deletion, where we randomly remove 10%, 15%, and 20% of words from the original text. Second, Synonym Substitution, replacing 10%, 15%, and 20% of words with their synonyms using WordNet (Miller, 1994). Lastly, we leverage the WikiMIA-paraphrased dataset (Zhang et al., 2024), which offers ChatGPT-generated rephrased versions of the original WikiMIA Shi et al. (2024a) texts while preserving their meaning.
We evaluate the effectiveness of baselines and Con-ReCall after transforming texts using the above techniques. Our experiments are conducted using Pythia-6.9B (Biderman et al., 2023) and LLaMA-30B (Touvron et al., 2023a) models on the WikiMIA-32 (Shi et al., 2024a) dataset. Table 3 presents the AUC performance for each method under various text manipulation scenarios. The results demonstrate that Con-ReCall consistently outperforms baseline methods across all text manipulation techniques, maintaining its superior performance even when faced with altered versions of the target data. This robustness underscores Con-ReCall’s effectiveness in real-world scenarios where data may undergo various transformations.
@@ 6 Conclusion
Acknowledgement
The work is supported by a National Science Foundation CAREER award #2339766, a research award NSF #2331966, University of California, Merced, University of Queensland, and the Ministry of Education, Singapore, under the Academic Research Fund Tier 1 (FY2023) (Grant A-8001996-00-00). The views and conclusions are those of the authors and should not reflect the official policy or position of the U.S. Government.